# Efficient Time Series Processing for Transformers and State-Space Models through Token Merging

### v1

@@ 5.1 Token merging in pretrained models
Regardless, token merging substantially increases the throughput of most models, up to , often with no change in forecasting quality as table 1 shows. In some experiments, token merging even improves the MSE. In line with the formal analysis of potential speed up from token merging conducted in section 3, we generally observe higher accelerations for larger models, as more subsequent layers can profit from already merged tokens.
@@
5.2 Token merging during training
Here, we apply token merging during training to reduce the model’s sensitivity to the algorithm at inference time. As shown in figure 2, models trained with token merging often outperform those trained without it, even if token merging is not applied during testing. This approach enables us to accelerate models such as Autoformer on the Traffic dataset without sacrificing accuracy, which was previously not feasible when applying token merging only during inference. Additionally, token merging accelerates the training process itself by up to for Autoformer on the Traffic dataset.
@@ 5.2 Token merging during training
Traffic |
5.3 Scaling to large models
Foundation models are getting more relevant across domains, including NLP [39], computer vision [40], and time series processing [22].
However, these models have high computational requirements. Thus, we investigate token merging for foundation models on Chronos, a univariate probabilistic model, in zero-shot forecasting setting [7]. We apply token merging during inference only, as training Chronos from scratch is not within the scope of this work.
In all our experiments, we find Pareto optimal points with token merging. For four out of five datasets token merging improves both accuracy and throughput simultaneously (see section A.3). Our results demonstrate that it is often beneficial to choose a larger Chronos model with token merging over a smaller one without as in figure 3.
We report our results in table 2, choosing the best Chronos model without token merging as reference. We illustrate two cases: 1) Selecting the token merging setting that provides the best MSE, 2) selecting the setting with the fastest throughput. For 2), we constrain the MSE of token merging trials to be lower than the second-best model without token merging. In addition, we allow a maximum increase in MSE of compared to the reference. In our experiments we can improve Chronos MSE by up to and speed up inference by .
5.4 Merging patterns
@@
We observe three distinct merging patterns when combining tokens in transformer architectures.
@@
Increasing MSE As the number of merged token increases the MSE increases almost monotonically (see figure 3). This behavior can be explained due to a loss of information when combining multiple tokens and also occurs in the vision domain [10].
@@
Constant MSE For the vanilla Transformer on ETTh1 and for FEDformer on ETTh1, Weather, Electricity, and Traffic, we observe a constant MSE when applying token merging as shown in figure 5. For the Transformer model, we find all tokens to be similar after the first attention block. Thus, token merging does not affect the model performance. Nevertheless, we find that in most cases, these models still provide reasonable forecasts. In our experiments, transformer models trained on larger or more complex datasets containing more variates do not show this behavior. We argue that this might be a limitation of transformers on small time series datasets [41, 42]. Still, token merging successfully improves the throughput while maintaining accuracy for these models.
@@
Decreasing MSE Token merging increases forecasting quality, most prominently in Chronos models as in figure 3. We explain this behavior in section 5.5.
5.5 When does token merging improve model performance
@@
In our experiment, applying token merging sometimes improves MSE. Our hypothesis is that averaging similar tokens smoothes the time series, reducing noise and acting as a low-pass filer. To validate our hypothesis, we low-pass filter the input time series using Gaussian kernels without token merging in figure 4. On ETTh1 and Traffic, both token merging and Gaussian filtering improve the MSE. On the Electricity dataset, token merging and Gaussian filtering do not positively impact the MSE. All of these observations are in line with our hypotheses. Applying token merging together with the Gaussian kernel leads to the best results. Other averaging kernels were significantly worse. We show additional results on ETTm1 and Weather in section A.4.
@@ 5.6 Dependencies on input length
Token merging effectively reduces the number of tokens in a transformer layer. Here, we explore if we can achieve similar accelerations while maintaining the same prediction quality by varying the number of input samples . For better comparison, we keep the predicted time series snippet fixed and only vary the input sequence.
Our results demonstrate that varying the input length cannot replace token merging (see also section A.5). In figure 6, we investigate input length dependence for two objectives in more detail: First, we explore the token merging setup that leads to the best MSE and compare the results to the model without merging. Here, token merging yields considerable throughput increases while improving predictive quality at the same time. Secondly, we compare the fastest model with token merging, which shows no quality decreases, to a standard model. We find models with token merging to scale favorable to long sequences compared to models without merging.
5.7 Redundancy of input tokens
First, we investigate whether scaling the number of input tokens increases average redundancy on the ETTh1 dataset. As demonstrated in figure 7(a), the same relative number of tokens are merged for a given merging threshold, independent of input length. Therefore, we suggest scaling the number of merged tokens in each layer linearly with the input length.
Positional embeddings add information about the location of a token within a sequence. As a result, two identical tokens without positional embeddings may show considerable differences when positional embeddings are included, potentially preventing merging. However, figure 7(a) shows that this effect on token merging is only marginal.
5.8 Dynamic merging
A fixed merging objective allows for batch processing without needing to pad individual time series to the same length. However, it enforces a fixed among independent batch elements, which might not always be optimal. Determining the number of tokens to be merged dynamically using a similarity threshold might increase quality as no dissimilar tokens are combined. Here, we leverage the single-sample case to explore dynamic merging in optimal conditions. From a practical perspective, this case might be relevant for on-device applications like smartphones or automated driving.
In figure 7(b), we compare token merging utilizing a fixed to dynamic merging varying the cosine similarity threshold. Dynamic merging improves quality slightly in most settings. Therefore, we suggest using a fixed merging schedule for batch applications and dynamic merging just for the single-sample case. There is no equivalent to dynamic merging schedules as they are similarity-based and strongly layer-dependent. We report FLOPs as we observe substantial execution overhead in time measurements.
5.9 Token merging in state-space models
Token merging | Accel. | Accuracy |
No merging | 1.00 | 78.9 % |
Local merging | 74.0 % | |
Global merging | 2.93 | 69.4 % |
We explore token merging in HyenaDNA. We use a classification task for the experiment, where the data consists of long genomic sequences with nucleotides each. Our local merging with featuring linear complexity and locality bias outperforms global merging with and quadratic complexity. Table 3 illustrates that local merging achieves substantially larger speed up and better accuracy than global merging. This experiment indicates that architecture and domain-specific biases are important when applying token merging to a task.
Local merging accelerates HyenaDNA with a decrease in accuracy. To the best of our knowledge, this is the first study that investigates merging individual states in state-space models to improve their sequence modeling performance.
@@ 6 Conclusion
In this work, we explore token merging in the time series domain for the first time. We conduct an extensive empirical study on transformer architectures and state-space models in diverse settings using various models and datasets. We demonstrate that token merging can successfully accelerate pretrained models and sometimes even improve their prediction quality. We further introduce a domain-specific local merging algorithm with variable complexity and illustrate its effectiveness on the Hyena model. On the long-range Dummy Mouse Enhancers Ensembl dataset, this method outperforms traditional token merging approaches in terms of throughput and accuracy.
Finally, we conduct several ablation studies to investigate when token merging is most effective, including sequence length, positional embedding, and single-sample inference settings.
We hope that token merging will have a positive effect on reducing the resource consumption and environmental impact of time series models.
Limitations
@@
In our work, we restrict the merging algorithm to the encoder part of all transformer architectures. Moreover, we divide all tokens into two sets and restrict merging to occur only between tokens from different sets. Future work can explore more flexible merging schemes for time series-specific architectures. Moreover, we do not conduct ablations on all possible hyperparameters due to the large number of architectures and datasets evaluated in this work. Lastly, we focus primarily on common metrics like MSE and accuracy and do not explore the impact of token merging on other factors, such as robustness to distribution shifts or adversarial attacks [43, 44].
Disclaimer
The results, opinions and conclusions expressed in this publication are not necessarily those of Volkswagen Aktiengesellschaft.
Appendix A Appendix
Supplementary material such as derivations, further details and additional results are listed below.
A.1 Derivations
In the following we derive our theoretical results in section 3.
@@ A.1 Derivations
Merging speed up bound
@@
We roughly estimate the upper bound of the speed up we can achieve by merging tokens in a -layer transformer model. Therefore, we only consider attention due to its quadratic scaling with . We disregard additional effects reducing speed up such as merging overhead to estimate the upper bound. Further, we assume merging half of the tokens each layer. The attention in the first layer is unaffected by merging as we apply token merging between the attention and MLP.
A.2 Experiments

### v3

@@ 5.1 Local merging in pretrained models
Regardless, our local merging in the encoder together with our casual merging in the decoder substantially increase the throughput of most models, up to , often with no change in forecasting quality, as table 1 shows. In some experiments, local merging even improves the MSE. In line with our formal analysis of potential speed-up from token merging in section 3, we generally observe higher accelerations for larger models, as more subsequent layers can profit from already merged tokens. Independent of model size, local merging finds Pareto optimal points in of settings and has no negative effect in the remaining cases.
@@
5.2 Local merging during training
Here, we apply local merging during training to reduce the models’ sensitivities to the algorithm at inference time. As shown in figure 2, models trained with token merging often outperform those trained without it, even if token merging is not applied during testing. This approach enables us to accelerate models such as Autoformer on the Traffic dataset without sacrificing accuracy, which was previously not feasible when applying token merging only during inference. Additionally, local merging accelerates the training process itself by up to for Autoformer on the Traffic dataset.
5.3 Scaling to large models
Foundation models are getting more relevant across domains, including NLP (Touvron et al., 2023), computer vision (Kirillov et al., 2023), and time series processing (Das et al., 2023).
However, these models have high computational requirements. Therefore, accelerating foundation models without the need for additional fine-tuning is especially important. Thus, we investigate local merging for foundation models on Chronos, a univariate probabilistic model, in a zero-shot forecasting setting (Ansari et al., 2024) and apply local merging only during inference.
In all our experiments, we find Pareto optimal points with token merging. For four out of five datasets, local merging improves both accuracy and throughput simultaneously (see section D.1). Our results demonstrate that it is often beneficial to choose a larger Chronos model with token merging over a smaller one without, as in figure 3.
We report our results in table 2, choosing the best model without token
merging as reference. We illustrate two cases: 1) Selecting the token merging setting that provides the best MSE, and 2) selecting the setting with the fastest throughput. For 2), we constrain the MSE of token merging trials to be lower than the second-best model without token merging. In addition, we allow a maximum increase in MSE of compared to the reference. In our experiments, we can improve Chronos MSE by up to and speed-up inference by .
@@ 5.3 Scaling to large models
| Traffic | |||||
5.4 Token merging in state-space models
We compare local and global token merging in HyenaDNA (Grešová et al., 2023) and Mamba (Gu & Dao, 2023), for two objectives: the largest speed-up and the best prediction quality. We use a classification task, where the data consists of long genomic sequences with nucleotides each. Our local merging with featuring linear complexity and locality bias outperforms global merging with and quadratic complexity. Table 3 illustrates that local merging achieves substantially larger speed-up and better accuracy than global merging. This experiment indicates that architecture and domain-specific biases are important when applying token merging.
Local merging accelerates HyenaDNA up to with a decrease in accuracy, whereas global merging substantially reduces the accuracy by . Utilizing less aggressive merging schemes, local merging even boosts accuracy by while still accelerating HyenaDNA . In Mamba models, local merging achieves even higher accelerations up to . In our experiments, the similarity computation of local merging adds of additional execution time to every Hyena block. For global merging, however, this is substantially higher with , highlighting the importance of local merging’s linear complexity. To our knowledge, this is the first study of merging individual states in state-space models to improve their sequence modeling performance.
| Token merging | Hyena | Mamba | ||
| Accel. | Accuracy | Accel. | Accuracy | |
| No merging | 1.00 | 78.9 % | 1.00 | 76.0 % |
| Local mergingfastest | 74.0 % | 74.4 % | ||
| Local mergingbest | 80.6 % | 76.0 % | ||
| Global mergingfastest | 2.93 | 69.4 % | 2.81 | 74.0 % |
| Global mergingbest | 1.15 | 80.2 % | 1.27 | 76.4 % |
5.5 Dynamic token merging
Dynamic token merging mitigates the issue of dissimilar tokens being merged, potentially improving quality.
Here, we leverage the single-sample case with batch size to explore dynamic merging in optimal conditions. From a practical perspective, this case might be relevant for on-device applications like smartphones or automated driving. We further utilize small batch sizes of elements as they might appear in real-world applications.
In figure 4, we compare token merging utilizing a fixed to dynamic merging, varying the cosine similarity threshold. Dynamic merging improves quality slightly in most settings.
With batch size , it is marginally worse compared to the optimal single-sample case.
Therefore, we suggest using a fixed merging schedule for applications with large batches and dynamic merging for cases with few batch elements. There is no equivalent to dynamic merging schedules as they are similarity-based and strongly layer-dependent. We report FLOPs as we observe substantial execution overhead in time measurements.
6 Further investigations
We first explore different local merging outcomes regarding acceleration and forecasting performance. Next, we find dataset- and model-specific properties explaining why, in some cases, local merging can improve forecasting quality. Lastly, we investigate if we can achieve same effects on model acceleration and forecasting performance with simpler methods than local merging. We further explore different token similarity measures in section E.1, compare merging with pruning in section E.2, and investigate the influence of tokenization on local merging in section E.3.
6.1 Merging outcomes
@@
We observe three distinct merging outcomes when combining tokens in transformer architectures.
@@
Increasing MSE As the number of merged tokens increases, the MSE increases almost monotonically (see figure 3(b)). This behavior can be explained due to a loss of information when combining multiple tokens and also occurs in the vision domain (Bolya et al., 2023).
@@
Constant MSE For the vanilla Transformer on ETTh1 and for FEDformer on ETTh1, Weather, Electricity, and Traffic, we observe a constant MSE when applying token merging as shown in figure 5. For the Transformer model, we find all tokens to be similar after the first attention block. Thus, token merging does not affect the model performance. Nevertheless, we find that in most cases, these models still provide reasonable forecasts. In our experiments, transformer models trained on larger or more complex datasets containing more variates do not show this behavior. We argue that this might be a limitation of transformers on small time series datasets (Zeng et al., 2023; Li et al., 2023). Still, token merging successfully improves the throughput while maintaining accuracy for these models.
@@
Decreasing MSE Token merging increases forecasting quality, most prominently in Chronos models, as in figure 3(a). We explain this behavior in section 6.2.
6.2 Improvement of forecasting quality
We explore if the potential benefits of token merging can be predicted through dataset- and model-specific properties without requiring downstream task evaluation.
@@ 6.2 Improvement of forecasting quality
Selective low-pass filter We hypothesize that local merging improves forecasting quality by selectively reducing noise. Averaging similar tokens smoothes the time series, acting as an adaptive low-pass filter. To validate our hypothesis, we low-pass filter the input time series using Gaussian kernels without token merging in figure 6. On ETTh1, both local merging and Gaussian filtering improve the MSE. On the Electricity dataset, token merging and Gaussian filtering do not positively impact the MSE. All of these observations are in line with our hypothesis. Applying token merging together with the Gaussian kernel leads to the best results. Other averaging kernels were significantly worse. We show additional results in section E.4.
Dataset properties We find properties of the target dataset that are particularly amenable to token merging. Using metrics from signal processing, we can predict how well local merging will perform on a new dataset prior to evaluation. Improvement in forecasting quality due to local merging in table 2 correlates with the spectral entropy of the dataset. Specifically, local merging achieves higher quality gains on high entropy datasets, such as ETTm1, ETTh1, and Traffic (see table 4). We argue that local merging removes unnecessary information from complex signals with high entropy using its selective smoothing ability. This allows the model to focus on only the relevant patterns of a signal and to achieve better prediction quality. Besides the spectral entropy, the same correlation is evident in the total harmonic distortion (THD). Local merging adaptively low-pass-filters noisy distorted signals to condense the most relevant patterns and effectively improves the signal-to-noise ratio. The greater noise in ETTm1, ETTh1, and Traffic compared to Electricity and Weather can also be visually inspected in the respective frequency spectrum in section E.5. Therefore, we expect a larger improvement of prediction quality when applying local merging on high entropy signals with a low signal-to-noise ratio. While less significant, local merging still improves efficiency on low noise datasets, as we discuss in section E.6 in detail.
| Dataset | MSEΔ | Spectral entropy | THD |
| ETTm1 | |||
| ETTh1 | |||
| Traffic | |||
| Electricity | |||
| Weather |
Model properties Across all datasets, we identify model-specific properties that benefit local merging. For this, we analyze the average cosine similarity of tokens in the models from table 1 after the first transformer layer. Local merging accelerates models, such as the Non-stationary Transformer, that learn more similar token representations without quality degradation. For models with dissimilar token representations, like the Informer, we observe quality degradations when applying local merging, as table 5 shows.
| Model and dataset | MSEΔ | Token similarity |
| Informer 2 Layers Traffic | ||
| Informer 4 Layers Electricity | ||
| Informer 8 Layers ETTh1 | ||
| Informer 6 Layers Weather | ||
| Informer 6 Layers ETTm1 | ||
| Non-stationary 10 Layers ETTh1 | ||
| Non-stationary 8 Layers ETTh1 | ||
| Non-stationary 6 Layers Weather | ||
| Transformer 10 Layers ETTm1 |
@@ 6.3 Dependencies on input length
Token merging effectively reduces the number of tokens in a transformer layer. Here, we explore if we can achieve similar accelerations while maintaining the same prediction quality by varying the number of input samples . For better comparison, we keep the predicted time series snippet fixed and only adjust the input sequence.
Our results demonstrate that varying the input length cannot replace local merging (see also section E.7). In figure 7, we investigate input length dependence for two objectives in more detail: First, we explore the token merging setup that leads to the best MSE and compare the results to the model without merging. Here, local merging yields considerable throughput increases while improving predictive quality at the same time. Second, we compare the fastest model with token merging, which shows no quality decreases, to a standard model. We find models with local merging to scale favorable to long sequences.
We further explore the redundancy of input tokens including the influence of the positional embedding in section E.6.
@@ 7 Conclusion
In this work, we explore token merging in the time series domain for the first time. We conduct an extensive empirical study on transformer architectures and state-space models in diverse settings using various models and datasets. We demonstrate that token merging can successfully accelerate pretrained models and sometimes even improve their prediction quality. We further introduce a domain-specific local merging algorithm with variable complexity and illustrate its effectiveness on Hyena and Mamba models. Additionally, local merging is the first causal token merging scheme, which we successfully apply in transformer decoders. Finally, we conduct several ablation studies to investigate when token merging is most effective, including spectral properties of the analyzed dataset and model- and algorithm-specific properties. We hope that token merging will have a positive effect on reducing the resource consumption and environmental impact of time series models.
@@
Limitations In our work, we divide all tokens into two sets and restrict merging to occur only between tokens from different sets. Future work can explore more flexible merging schemes for time-series-specific architectures. Moreover, we do not conduct ablations on all possible hyperparameters due to the large number of architectures and datasets evaluated in this work.
Impact Statement
We demonstrate large accelerations and considerable quality gains throughout a broad range of time series architectures. Our local merging can improve training efficiency and accelerate already trained models without any fine-tuning. This is especially important for emerging foundation models, which require considerable computational resources. We hope that local merging will contribute to more sustainable time series models and reduce their environmental impact.
Appendix A Related work
Here, we discuss related work in greater detail.
Time series transformers In recent years, many transformer architectures with inductive biases for time series have been proposed, successfully outperforming classical and other deep-learning-based methods in time series forecasting quality like recurrent neural networks (Li et al., 2019). Most of them focus on reducing complexity by modifying the attention mechanism. LogTrans uses LogSparse attention (Li et al., 2019), while Informer focuses only on the most relevant queries using ProbSparse attention (Zhou et al., 2021). Additionally, many architectures adopt decomposition techniques to model trend and seasonal patterns (Woo et al., 2022; Wu et al., 2021; Zhou et al., 2022; Liu et al., 2022b). Autoformer leverages autocorrelation as a sequence-based similarity measure in the attention mechanism (Wu et al., 2021). FEDformer uses the frequency domain to model time series effectively (Zhou et al., 2022). Non-stationary Transformers further mitigate the effect of the time series distribution changing over time (Liu et al., 2022b). Other works apply hierarchical attention (Liu et al., 2022a; Cirstea et al., 2022), embed subsequences as tokens to capture local semantic information (Nie et al., 2023), or leverage attention between time series variates to better model multivariate patterns (Zhang & Yan, 2023; Liu et al., 2023). Due to their success in the vision and NLP domain, transformer-based foundation models have lately emerged for time series, often used in zero-shot settings. Many works focus on training transformers directly on large and diverse time series datasets, usually with billions of tokens (Garza & Mergenthaler-Canseco, 2023; Das et al., 2023; Rasul et al., 2023; Woo et al., 2024). Inspired by the success of foundation models in NLP, the recently proposed Chronos model converts continuous time series data into a fixed vocabulary and is trained on both real-world and synthetic data (Ansari et al., 2024). Besides, other research branches focus on fine-tuning vision or NLP models for time series (Zhou et al., 2023) and on applying large language models directly on time series data (Gruver et al., 2023).
State-space models Due to the quadratic scaling of the attention mechanism, transformer architectures suffer from significant computational cost when processing very long sequences. Recently, state-space models have shown promising results in overcoming the quadratic complexity of transformers with respect to input length. Linear state-space layers solve the sequential processing requirement of RNNs through linear state-space representations (Gu et al., 2021). The S4 model reduces memory requirements by conditioning the state-space matrix with a low-rank correction (Gu et al., 2022). By using implicit convolutions and a data-aware gating mechanism, Hyena (Poli et al., 2023) became one of the first state-space model architectures to match transformers on NLP tasks. Later work uses hardware-aware algorithms to improve the performance of state-space models on modern accelerators (Gu & Dao, 2023).
Reducing tokens Many works reduce the number of processed tokens to increase the efficiency of transformers in computer vision and NLP, often by pruning (Meng et al., 2022; Goyal et al., 2020). Marin et al. (2021) merge tokens in ViT architectures to reduce the loss of information associated with pruning. Bolya et al. (2023) enhance the token merging algorithm, which they successfully apply to already trained encoder-only models. Besides initial work on classification tasks (Bolya et al., 2023), subsequent work applies token merging to diffusion models (Bolya & Hoffman, 2023). Kim et al. (2024) combine merging and pruning, while other works investigate optimal merging and pruning rates (Bonnaerens & Dambre, 2023; Chen et al., 2023). Concurrent work adapts token merging to preserve the spectral properties of the token space (Tran et al., 2024). However, their merging algorithm still has quadratic complexity, making it unsuitable for long sequence processing.
Sparse attention and token skipping Besides reducing the number of tokens, sparse attention (Child et al., 2019; Li et al., 2019; Zhou et al., 2021; Wu et al., 2021) and token skipping (Raposo et al., 2024) also decrease the computational requirements of transformer models. Sparse attention computes a subset of the attention matrix. Therefore, it can only accelerate the attention mechanism itself and not the subsequent MLP, in contrast to reducing the number of tokens during token merging. According to Marin et al. (2021), this MLP can take over of the total computation in a ViT layer. Further, altering the network architecture from full attention to sparse attention might require a retraining of the model. Concurrent work, such as token skipping (Raposo et al., 2024), involves the selection of a subset of tokens to be processed in a transformer layer. However, it has only been shown in NLP when training from scratch. In contrast to sparse attention and token skipping, token merging can accelerate already trained models and does not require any training data or fine-tuning. This is especially important for recent foundation models, which are expensive to train. In our experiments in sections 5.1 and 5.2, token merging successfully accelerates Informer and Autoformer, which already employ sparse attention. We therefore consider token merging as an orthogonal approach.
Here, we propose the first token merging algorithm for the time series domain, which extends beyond previous investigations of token merging in ViTs (Bolya et al., 2023; Bolya & Hoffman, 2023). We systematically evaluate the potential to reduce computational effort in time-series-specific transformer architectures and state-space models.
Appendix B Local token merging for time series
In the following, we provide derivations and more details on our local merging algorithm. We further discuss the interplay of time-series-specific inductive biases and our token merging method.
B.1 Derivations
Here, we derive our theoretical results in section 3.
@@ B.1 Derivations
Merging speed-up bound
@@
We roughly estimate the upper bound of the speed-up we can achieve by merging tokens in a -layer transformer model. Therefore, we only consider attention due to its quadratic scaling with . We disregard additional effects reducing speed-up, such as merging overhead, to estimate the upper bound. Further, we assume merging half of the tokens in each layer. The attention in the first layer is unaffected by merging, as we apply token merging between the attention and MLP.
B.2 Inductive biases
Time series feature domain-specific characteristics, including temporal causality, long sequences, periodicity, trends, and sparsity in the frequency domain. Here, we discuss the interplay of these properties with our token merging algorithm.
Our local merging algorithm is specifically designed to exploit two core properties of time series: First, it preserves temporal causality, as real-world time series are generated by causal processes. Second, it maintains linear complexity as time series often consist of very long token sequences (Godahewa et al., 2021; Grešová et al., 2023). This way, we design a very universal token merging scheme, applicable to many model architectures and datasets, as we show in our experiments.
We conduct new investigations where we trace the tokens being merged throughout the transformer model and show that token merging can exploit periodicity and trends without explicitly modeling these inductive biases. As illustrated in figure 8(b), our global merging for time series combines local and global information. However, we did not implement these properties as hard inductive biases to maintain the universality of our algorithm: This way, token merging also performs well on sequential data that does not exhibit trend or periodicity, such as DNA sequences (Grešová et al., 2023), as we show in section 5.4. Stock prices typically also do not have regular periodic patterns. Further, introducing a periodic bias to the neighborhood of our local merging algorithm would break causality, making it inapplicable to decoders.
Autoformer and FEDformer transform the tokens to the frequency space. Autoformer specifically focuses on the autocorrelation. Here, our token merging natively exploits sparsity in the frequency domain and autocorrelation space.
Our token merging algorithm can exploit inductive biases for time series, including periodicity, trends, and sparsity in the frequency or autocorrelation space, but it is not limited to those. This way, it is universally applicable to many architectures and datasets. Further, it features causality and low complexity as inductive biases for 1d-sequence processing.
Appendix C Experiments